y = -x*2/8 + 11x/8 + 3/2
Find the equation of the axis of symmetry, explaining your method. Use this information to find the coordinates of the vertex, giving your answers to one decimal place.

Answer by Alan3354(69443)

y = -x*2/8 + 11x/8 + 3/2
Find the equation of the axis of symmetry, explaining your method. Use this information to find the coordinates of the vertex, giving your answers to one decimal place.
--------------
y = -x^2/8 + 11x/8 + 3/2 ---- Use ^ (Shift 6) for exponents.
* --> multiply
==========================
f(x) = -x^2/8 + 11x/8 + 3/2
-----
AOS is x = -b/2a = (-11/8)/(-1/4)
x = 11/2
-------
Vertex is (11/2,f(11/2))
f(11/2) = -121/32 + 121/16 + 3/2 = 169/32
Vertex (11/2,169/32)